OpenStudy (anonymous):

determine the exact value for sine and cosine of 225°

OpenStudy (demonchild99):

hmmm what grade r u in???

OpenStudy (anonymous):

11 why

OpenStudy (demonchild99):

hmmmm....give me a minute...im only in ninth

OpenStudy (anonymous):

okay

OpenStudy (demonchild99):

brb

OpenStudy (anonymous):

\[\sin (225^{0}) = \sin(180^{0} +45^{0})\] \[\cos(225^{0}) = \cos180^{0} +45^{0})\]

OpenStudy (anonymous):

use the formula of \[\sin(180^{0}+\theta) = - \sin \theta \] \[\cos(180^{0}+\theta) = - \cos\theta \]

OpenStudy (demonchild99):

??

OpenStudy (demonchild99):

yeah sorry dude...im totally lost...id love to help but i just dont know how to do it

OpenStudy (anonymous):

okay

OpenStudy (demonchild99):

sorry